# Reentrant Superconductivity in Zeeman Fields

### v1

@@ Abstract
and $\bm{H}=H\bm{e}_{x}$, where both the SOI and the pair potential have two components. The anomalous Green’s function has complicated structure as shown in Eq. (19). The three induced pairing correlations in Table 1 characterize the stability and the instability of the mixed superconducting state. The phase diagram is shown in Fig. 1(c), where we choose $\bm{g}=(g,0,g)$ in Eq. (24), fix $\alpha_{x}$ at 0.75 $T_{0}$, and determine the phase boundary from the minima of the free-energy. At $H=0$, odd-frequency spin-triplet Cooper pairing correlation proportional to $\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}$ make the superconducting states unstable. Both $\alpha_{x}$ and $\alpha_{y}$ suppress the pair potential $d_{z}$, whereas only $\alpha_{y}$ suppresses $d_{x}$. As a result, single-component superconducting states with $\bm{d}=(d_{x},0,0)$ are realized for small $H$. With increasing Zeeman fields from zero, the transition temperature first decreases because odd-frequency spin-singlet pairing correlation proportional to $\bm{d}\cdot\bm{H}$ suppresses such superconducting state. The results for $\alpha_{y}=1.25T_{0}$ show that superconductivity vanishes for $0.55T_{0}<\mu_{\mathrm{B}}H<1.12T_{0}$ even at $T=0$. However, another single component superconductivity with $\bm{d}=(0,0,d_{z})$ appears for high Zeeman fields $1.12T_{0}<\mu_{\mathrm{B}}H$. A Zeeman field does not affect $d_{z}$ component because of $\bm{d}\cdot\bm{H}=0$. In addition, the spin-singlet pairing correlation proportional to $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ stabilize superconducting states because it increases the superfluid density. We conclude that spin-singlet Cooper pairs induced by the interplay between the SOI and a Zeeman field cause the reentrant superconductivity. A spin-singlet Cooper pair in high Zeeman fields is an object which does not align with institution. However, our calculation clearly indicates its existence. A recent experiment on UTe2 shows a sign of spin-singlet pairs in the high field phase [13].
Figure 2: The critical magnetic field $H_{c}$ is plotted as a function of temperature for a single-band spin-triplet superconductors with Rashba SOI. The pair potential is chosen as Eq. (16a) in (a) and Eq. (16b) in (b).
Finally, we demonstrate the reentrant superconductivity in a usual single-band SC in two-dimension,
$\check{H}_{\mathrm{N}}(\bm{k})=$ $\left(\frac{\bm{k}^{2}}{2m}-\mu\right)\hat{\sigma}_{0}+\bm{\alpha}_{\bm{k}}\cdot\hat{\bm{\sigma}}+\mu_{\mathrm{B}}{H}\hat{\sigma}_{x},$ (13)
$\hat{\Delta}(\bm{k})=$ $i\bm{d}_{\bm{k}}\cdot\hat{\bm{\sigma}}\,i\hat{\sigma}_{y},$ (14)
$\bm{\alpha}_{\bm{k}}=$ $\alpha\left(\bar{k}_{y}\bm{e}_{x}-\bar{k}_{x}\bm{e}_{y}\right),$ (15)
where we consider the Rashba SOI and $\bar{k}_{j}=k_{j}/k_{F}$ for $j=x,y$. We consider the two pair potentials as
$\bm{d}_{\bm{k}}=$ $d_{x}\bar{k}_{x}\bm{e}_{x}+d_{z}\bar{k}_{y}\bm{e}_{z},$ (16a)
$\bm{d}_{\bm{k}}=$ $d_{x}\bar{k}_{y}\bm{e}_{x}+d_{z}\bar{k}_{x}\bm{e}_{z}.\$(16b)
We assume that the attractive interaction at $p$-wave channel in the gap equation. The phase diagrams for Eqs. (16a) and (16b) are shown in Fig. 2(a) and (b), respectively. The results for $\alpha=1.5T_{0}$ in (a) show ususal single superconducting phase, whereas those in (b) show two separated superconducting phases. The phase diagram depends sensitively on the relative $\bm{k}$ dependence between $\bm{\alpha}_{\bm{k}}$ and $\bm{d}_{\bm{k}}$ for large $\alpha$. Such tendency seems to be weaker for smaller $\alpha$. The reentrant superconductivity can be seen in the results for $\alpha=1.2T_{0}$ and $\alpha=T_{0}$ in both Figs. 2(a) and (b).
@@
Conclusion: We have studied the stability of a spin-triplet superconducting state described by a $\bm{d}$ vector in the presence of a spin-orbit interaction $\bm{\alpha}$ and a Zeeman field $\bm{H}$. The spin-orbit interaction for $\bm{\alpha}\times\bm{d}\neq 0$ and Zeeman fields for $\bm{H}\cdot\bm{d}\neq 0$ make the superconducting state unstable and decrease the transition temperature. The instability is characterized well by the appearance of odd-frequency Cooper pairs because they decrease the superfluid density. We demonstrate that a spin-triplet superconducting state exhibits the Zeeman-field-induced superconductivity when a scaler chiral product of $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ remains finite. Under the condition, even-frequency spin-singlet even-parity Cooper pairs stabilize the spin-triplet superconductivity at high Zeeman fields. Based on the obtained results, we propose a Hamiltonian of superconducting states which indicate the reentrant superconductivity in Zeeman fields. Although the Jaccarino-Peter compensation has been a reasonable story for the Zeeman-field-induced superconductivity, our conclusion provides an alternative physical picture for the phenomenon. In addition, our theory can be applied also to superconductors preserving time-reversal symmetry in the absence of a magnetic field.
@@
Acknowledgments: T. S. is supported by JST SPRING, Grant Number JPMJSP2119. S. I. is supported by a Grant-in-Aid for Early-Career Scientists (JSPS KAKENHI Grant No. JP24K17010).
Appendix A Solusion of Gor’kov equation
@@
$H_{\mathrm{BdG}}(\bm{k})$ $=\begin{bmatrix}\xi_{-}+\bm{\alpha}\cdot\hat{\bm{\sigma}}+\bm{B}\cdot\hat{\bm{\sigma}}&0&0&i\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\\ 0&\xi_{+}-\bm{\alpha}\cdot\hat{\bm{\sigma}}+\bm{B}\cdot\hat{\bm{\sigma}}&-i\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}&0\\ 0&-i\bm{d}\cdot\hat{\bm{\sigma}}^{\ast}\,\hat{\sigma}_{y}&-\xi_{+}-\bm{\alpha}\cdot\hat{\bm{\sigma}}^{*}-\bm{B}\cdot\hat{\bm{\sigma}}^{*}&0\\ i\bm{d}\cdot\hat{\bm{\sigma}}^{\ast}\,\hat{\sigma}_{y}\ &0&0&-\xi_{-}+\bm{\alpha}\cdot\hat{\bm{\sigma}}^{*}-\bm{B}\cdot\hat{\bm{\sigma}}^{*}\end{bmatrix}$, (17)
$\xi_{\pm}$ $=\frac{1}{2m}(\bm{k}\pm\bm{K})^{2}-\mu,\quad\bm{B}=\mu_{\mathrm{B}}\bm{H},$ (18)
@@
$\check{F}(\bm{k},\omega_{n})=$ $-\bm{d}^{2}\,\check{Z}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\alpha^{2}+\bm{B}^{2})\,\bm{d}\cdot\hat{\bm{\sigma}}-2(\bm{B}\cdot\bm{d})\bm{B}\cdot\hat{\bm{\sigma}}+2(\bm{\alpha}\cdot\bm{d})\bm{\alpha}\cdot\hat{\bm{\sigma}}\right$.
$-2\hat{\xi}_{\bm{k}}\,\bm{\alpha}\cdot\bm{d}\hat{\rho}_{z}+2\,i\,\hat{\xi}_{\bm{k}}\,(\bm{B}\times\bm{d})\cdot\hat{\bm{\sigma}}$
$\left.+2i\,\omega_{n}\,\bm{B}\cdot\bm{d}+2\,\omega_{n}\,(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\,\hat{\rho}_{z}-2i\,\bm{\alpha}\times\bm{B}\cdot\,\bm{d}\,\hat{\rho}_{z}\right]\,i\hat{\sigma}_{y}\,(i\hat{\rho}_{y}),$ (19)
$\check{Z}(\bm{k},\omega_{n})=$ $\left\{\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\bm{\alpha}^{2}+\bm{B}^{2}\right\}^{2}\bm{d}^{2}-4(\hat{\xi}_{\bm{k}}\bm{B}-i\omega_{n}\bm{\alpha}\hat{\rho}_{z})^{2}\bm{d}^{2}+4(\bm{\alpha}\cdot\bm{d})^{2}(\bm{B}^{2}+\bm{d}^{2})+4(\bm{B}\cdot\bm{d})^{2}(\bm{\alpha}^{2}-\bm{d}^{2})$
$+4(\bm{\alpha}\times\bm{B}\cdot\bm{d})^{2}-8(\bm{\alpha}\cdot\bm{d})(\bm{B}\cdot\bm{d})(\bm{\alpha}\cdot\bm{B}),$ (20)
$=$ $\hat{\rho}_{x}\,\check{Z}(-\bm{k},-\omega_{n})\,\hat{\rho}_{x}.$ (21)
@@
The anomalous Green’s function is antisymmetric under the operation of interchanging two electron,
$\check{F}(\bm{k},\omega_{n})=-\hat{\rho}_{x}\,\check{F}^{\mathrm{T}}(-\bm{k},-\omega_{n})\,\hat{\rho}_{x},$ (22)
@@
where $\mathrm{T}$ represents the transpose of the Pauli matrices for spin meaning the exchange of two spins and $\hat{\rho}_{x}\cdots\hat{\rho}_{x}$ represents the exchange of the two valleys. Eq. (21) indicates that $\check{Z}$ is symmetric under such operation. The three components on the first line in Eq. (19) represent spin-triplet odd-valley-parity symmetry class and are linked to the pair potential through the gap equation,
@@
$i\left(\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\right)_{\alpha,\beta}=$ $-T\sum_{\omega_{n}}\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}\sum_{\gamma,\delta}g_{\alpha\,\beta;\gamma\,\delta}\hat{F}^{(y)}_{\gamma,\delta}(\bm{k},\omega_{n}),$ (23)
$g_{\alpha\,\beta;\gamma\,\delta}=$ $\sum_{\nu=x,y,z}g_{\nu}\left(i\hat{\sigma}_{\nu}\,\hat{\sigma}_{y}\right)_{\alpha,\beta}\;\left(i\hat{\sigma}_{\nu}\,\hat{\sigma}_{y}\right)^{\ast}_{\gamma,\delta},$ (24)
@@
where $\hat{F}^{(y)}$ is the $\hat{\rho}_{y}$ component of the anomalous Green’s function and $\bm{g}=(g_{x},g_{y},g_{z})$ represents the attractive interaction between two electrons at an $s$-wave channel. To draw Fig. 2, we assume an attractive interaction at $p$-wave channel as usual
$g_{\nu}(\bm{k}-\bm{k}^{\prime})=g_{\nu}\cos(\theta-\theta^{\prime}),\;\cos\theta=k_{x}/k_{F},\;\sin\theta=k_{y}/k_{F}.$ (25)
@@
To our knowledge, the two components proportional to $\hat{\xi}_{\bm{k}}$ at the second line in Eq. (19) do not play any important role in stabilizing superconducting states. Namely, the presence of these components does not change $T_{c}$ at all. The first two components at the last line belong to the odd-frequency symmetry class and make the superconducting state unstable. [3]. A Zeeman field parallel to $\bm{d}$ and a SOI perpendicular to $\bm{d}$ generate such odd-frequency pairs from the pair potential. As a result, these components decrease $T_{c}$. The last component at the third line in Eq. (19) belongs to even-frequency spin-singlet even-valley-parity symmetry class. This component stabilizes a superconducting state at high Zeeman field. However, it appears only when the three vectors have a finite spin chiral product $\bm{\alpha}\times\bm{H}\cdot\,\bm{d}$.
@@
$\check{F}_{\parallel}(\bm{k},\omega_{n})=$ $-\frac{1}{2}\left[(\hat{X}_{+}^{-1}+\hat{X}_{-}^{-1})\bm{d}\cdot\hat{\bm{\sigma}}-(\hat{X}_{+}^{-1}-\hat{X}_{-}^{-1})d\right]i\hat{\sigma}_{y}\,(-i)\hat{\rho}_{y},$ (26)
$\hat{X}_{\pm}$= $(\omega_{n}\pm i\mu_{\mathrm{B}}H)^{2}+(\hat{\xi}_{\bm{k}}\mp\alpha\,\hat{\rho}_{z})+d^{2}.$ (27)
@@
$\check{F}_{\perp}(\bm{k},\omega_{n})=$ $-\check{Z}_{\perp}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}-\alpha^{2}+\mu_{\mathrm{B}}^{2}H^{2})\bm{d}\cdot\hat{\bm{\sigma}}-2i\hat{\xi}_{\bm{k}}\,\mu_{\mathrm{B}}\,\bm{d}\times\bm{H}\cdot\hat{\bm{\sigma}}\right$.
$\left.+2\omega_{n}\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}+2i\mu_{\mathrm{B}}\bm{\alpha}\times\bm{d}\cdot\bm{H}\hat{\rho}_{z}\right]i\hat{\sigma}_{y}(-i)\hat{\rho}_{y},$ (28)
$\check{Z}_{\perp}=$ $\hat{\xi}_{\bm{k}}^{4}+2\hat{\xi}_{\bm{k}}^{2}(\omega_{n}^{2}-\alpha^{2}-\mu_{\mathrm{B}}^{2}H^{2})+(\omega_{n}^{2}+\alpha^{2}+\mu_{\mathrm{B}}^{2}H^{2})^{2}.$ (29)

### v3

@@ Abstract
and $\bm{H}=H\bm{e}_{x}$, where both the SOI and the pair potential have two components. Based on HTH-T phase diagram in Figs. 1(a) and (b) and the interpretation of the results by using odd-frequency Cooper pairing correlations, it is possible to predict the reentrant superconductivity in the mixed configuration in Eq. (14). At $H=0$, the odd-frequency pairing correlation proportional to $\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}$ make the superconducting states unstable. Both $\alpha_{x}$ and $\alpha_{y}$ suppress the pair potential $d_{z}$, whereas only $\alpha_{y}$ suppresses $d_{x}$. As a result, single-component superconducting states with $\bm{d}=(d_{x},0,0)$ would be realized for small $H$. In high Zeeman fields, the pair potential $d_{x}$ vanishes due to the odd-frequency correlation proportional to $\bm{d}\cdot\bm{H}$ and the superconductivity due to $d_{z}$ would be stabilized by the spin singlet pairing correlation proportional to $\bm{\alpha}\times\bm{d}\cdot\bm{H}$. The phase diagram obtained from the minima of the free-energy shown in Fig. 1(c) are consistent with the predictions, where we choose $\bm{g}=(g,0,g)$ in Eq. (S11) and fix $\alpha_{x}$ at 0.75 $T_{0}$. With increasing Zeeman fields from zero, the transition temperature first decreases for both $\alpha_{y}/T_{0}=1.0$ and 1.25. The results for $\alpha_{y}=1.25T_{0}$ show that superconductivity vanishes for $0.55T_{0}<\mu_{\mathrm{B}}H<1.12T_{0}$ even at $T=0$. Another single component superconductivity with $\bm{d}=(0,0,d_{z})$ appears for high Zeeman fields $1.12T_{0}<\mu_{\mathrm{B}}H$. For $\alpha_{y}=T_{0}$, the low-field phase and high-field phase are connected at the vending point. However, the $\bm{d}$ vector has only one component in the two phases: $d_{x}$ at low-field phase and $d_{z}$ at high-field phase. We have mathematically confirmed that the multi-component superconducting phase is absent because of the idealistic model setting in this paper. The conditions for the multi-component superconductivity will be discussed elsewhere. In Fig. S1 in SM, we demonstrate that the reentrant superconductivity happens even in usual single-band superconductors with the odd-parity potentials $\bm{d}_{\bm{k}}=-\bm{d}_{-\bm{k}}$ and $\bm{\alpha}_{\bm{k}}=-\bm{\alpha}_{-\bm{k}}$. These results indicate that the magnetic configuration in spin space $\bm{\alpha}\times\bm{d}\cdot\bm{H}\neq 0$ is essential for the ZFIS and that the physical interpretation of the phenomenon using odd-frequency pairs is valid.
We conclude that spin-singlet Cooper pairs induced by the interplay between the SOI and a Zeeman field cause the reentrant superconductivity. A spin-singlet Cooper pair in high Zeeman fields is an object which does not align with institution. However, our calculation clearly indicates its existence. A recent experiment on UTe2 shows a sign of spin-singlet pairs in the high field phase [18].
@@
Conclusion: We have studied the stability of a spin-triplet superconducting state described by a $\bm{d}$ vector in the presence of a spin-orbit interaction $\bm{\alpha}$ and a Zeeman field $\bm{H}$. The spin-orbit interaction for $\bm{\alpha}\times\bm{d}\neq 0$ and Zeeman fields for $\bm{V}\cdot\bm{d}\neq 0$ make the superconducting state unstable and decrease the transition temperature. The instability is characterized well by the appearance of odd-frequency Cooper pairs because they decrease the superfluid density. We demonstrate that a spin-triplet superconducting state exhibits the Zeeman-field-induced superconductivity when a scalar chiral product of $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ remains finite. Under the condition, even-frequency spin-singlet even-parity Cooper pairs stabilize the spin-triplet superconductivity at high Zeeman fields. Based on the obtained results, we propose a Hamiltonian of superconducting states which indicate the reentrant superconductivity in Zeeman fields. Although the Jaccarino-Peter compensation has been a reasonable story for the Zeeman-field-induced superconductivity, our conclusion provides an alternative physical picture for the phenomenon. In addition, our theory can be applied also to superconductors preserving time-reversal symmetry in the absence of a magnetic field.
@@
Acknowledgments: T. S. is supported by JST SPRING, Grant Number JPMJSP2119. S. I. is supported by a Grant-in-Aid for Early-Career Scientists (JSPS KAKENHI Grant No. JP24K17010). Y. A. is supported by a Grant-in-Aid for Scientific Research (JSPS KAKENHI Grant No. JP26K0692).
Supplemental Material for “Reentrant Superconductivity in Zeeman Fields”
Tomoya Sano, Kota Tabata, Satoshi Ikegaya, and Yasuhiro Asano
Department of Applied Physics, Hokkaido University, Sapporo 060-8628, Japan
I Solution of Gor’kov equation
@@
$H_{\mathrm{BdG}}(\bm{k})$ $=\begin{bmatrix}\xi_{-}+\bm{\alpha}\cdot\hat{\bm{\sigma}}+\bm{V}\cdot\hat{\bm{\sigma}}&0&0&i\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\\ 0&\xi_{+}-\bm{\alpha}\cdot\hat{\bm{\sigma}}+\bm{V}\cdot\hat{\bm{\sigma}}&-i\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}&0\\ 0&-i\bm{d}\cdot\hat{\bm{\sigma}}^{\ast}\,\hat{\sigma}_{y}&-\xi_{+}-\bm{\alpha}\cdot\hat{\bm{\sigma}}^{*}-\bm{V}\cdot\hat{\bm{\sigma}}^{*}&0\\ i\bm{d}\cdot\hat{\bm{\sigma}}^{\ast}\,\hat{\sigma}_{y}\ &0&0&-\xi_{-}+\bm{\alpha}\cdot\hat{\bm{\sigma}}^{*}-\bm{V}\cdot\hat{\bm{\sigma}}^{*}\end{bmatrix}$, (S1)
$\xi_{\pm}$ $=\frac{1}{2m}(\bm{k}\pm\bm{K})^{2}-\mu,\quad\bm{V}=\mu_{\mathrm{B}}\bm{H},$ (S2)
@@
$\check{G}(\bm{k},\omega_{n})=$ $\check{Z}^{-1}\left[\hat{z}_{D}(i\omega_{n}-\hat{\xi}_{\bm{k}})-d^{2}(i\omega_{n}+\hat{\xi}_{\bm{k}})-2(\bm{\alpha}\cdot\bm{d},\hat{\rho}_{z}+\bm{V}\cdot\bm{d})\,\bm{d}\cdot\hat{\bm{\sigma}}\right$.
$+\left.(\hat{z}_{D}+\bm{d}^{2})\bm{\alpha}\cdot\hat{\bm{\sigma}}\,\hat{\rho}_{z}-(\hat{z}_{D}-\bm{d}^{2})\bm{V}\cdot\hat{\bm{\sigma}}\right$], (S3)
$\check{F}(\bm{k},\omega_{n})=$ $\check{Z}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\bm{\alpha}^{2}+\bm{V}^{2})\,\bm{d}\cdot\hat{\bm{\sigma}}-2(\bm{V}\cdot\bm{d})\bm{V}\cdot\hat{\bm{\sigma}}+2(\bm{\alpha}\cdot\bm{d})\bm{\alpha}\cdot\hat{\bm{\sigma}}\right$.
$-2\hat{\xi}_{\bm{k}}\,\bm{\alpha}\cdot\bm{d}\hat{\rho}_{z}+2\,i\,\hat{\xi}_{\bm{k}}\,(\bm{V}\times\bm{d})\cdot\hat{\bm{\sigma}}$
$\left.+2i\,\omega_{n}\,\bm{V}\cdot\bm{d}+2\,\omega_{n}\,(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\,\hat{\rho}_{z}-2i\,\bm{\alpha}\times\bm{V}\cdot\,\bm{d}\,\hat{\rho}_{z}\right]\,\hat{\sigma}_{y}\,\hat{\rho}_{y},$ (S4)
@@
$-\undertilde{\check{F}}(\bm{k},\omega_{n})=$ $\hat{\sigma}_{y}\,\hat{\rho}_{y}\check{Z}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\bm{\alpha}^{2}+\bm{V}^{2})\,\bm{d}\cdot\hat{\bm{\sigma}}-2(\bm{V}\cdot\bm{d})\bm{V}\cdot\hat{\bm{\sigma}}+2(\bm{\alpha}\cdot\bm{d})\bm{\alpha}\cdot\hat{\bm{\sigma}}\right$.
$-2\hat{\xi}_{\bm{k}}\,\bm{\alpha}\cdot\bm{d}\hat{\rho}_{z}-2\,i\,\hat{\xi}_{\bm{k}}\,(\bm{V}\times\bm{d})\cdot\hat{\bm{\sigma}}$
$\left.+2i\,\omega_{n}\,\bm{V}\cdot\bm{d}-2\,\omega_{n}\,(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\,\hat{\rho}_{z}+2i\,\bm{\alpha}\times\bm{V}\cdot\,\bm{d}\,\hat{\rho}_{z}\right],$ (S5)
$\check{Z}(\bm{k},\omega_{n})=$ $\left\{\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}+\bm{d}^{2}-\bm{\alpha}^{2}+\bm{V}^{2}\right\}^{2}\bm{d}^{2}-4(\hat{\xi}_{\bm{k}}\bm{V}-i\omega_{n}\bm{\alpha}\hat{\rho}_{z})^{2}\bm{d}^{2}+4(\bm{\alpha}\cdot\bm{d})^{2}(\bm{V}^{2}+\bm{d}^{2})+4(\bm{V}\cdot\bm{d})^{2}(\bm{\alpha}^{2}-\bm{d}^{2})$
$+4(\bm{\alpha}\times\bm{V}\cdot\bm{d})^{2}-8(\bm{\alpha}\cdot\bm{d})(\bm{V}\cdot\bm{d})(\bm{\alpha}\cdot\bm{V}),$ (S6)
$=$ $\hat{\rho}_{x}\,\check{Z}(-\bm{k},-\omega_{n})\,\hat{\rho}_{x},$ (S7)
$\hat{z}_{D}=$ $(i\omega_{n}+\hat{\xi}_{\bm{k}})^{2}-(\bm{\alpha}\hat{\rho}_{z}+\bm{V})^{2}.$ (S8)
@@
The anomalous Green’s function is antisymmetric under the operation of interchanging two electrons,
$\check{F}(\bm{k},\omega_{n})=-\hat{\rho}_{x}\,\check{F}^{\mathrm{T}}(-\bm{k},-\omega_{n})\,\hat{\rho}_{x},$ (S9)
@@
where $\mathrm{T}$ represents the transpose of the Pauli matrices for spin meaning the exchange of two spins and $\hat{\rho}_{x}\cdots\hat{\rho}_{x}$ represents the exchange of the two valleys. Eq. (S7) indicates that $\check{Z}$ is symmetric under such operation. The three pairing correlations on the first line in Eq. (S4) belong to spin-triplet odd-valley-parity symmetry class and are linked to the pair potential through the gap equation,
@@
$i\left(\bm{d}\cdot\hat{\bm{\sigma}}\,\hat{\sigma}_{y}\right)_{\alpha,\beta}=$ $-T\sum_{\omega_{n}}\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}\sum_{\gamma,\delta}g_{\alpha\,\beta;\gamma\,\delta}\hat{F}^{(y)}_{\gamma,\delta}(\bm{k},\omega_{n}),$ (S10)
$g_{\alpha\,\beta;\gamma\,\delta}=$ $\sum_{\nu=x,y,z}g_{\nu}\left(i\hat{\sigma}_{\nu}\,\hat{\sigma}_{y}\right)_{\alpha,\beta}\;\left(i\hat{\sigma}_{\nu}\,\hat{\sigma}_{y}\right)^{\ast}_{\gamma,\delta},$ (S11)
@@
where $\hat{F}^{(y)}$ is the $\hat{\rho}_{y}$ component of the anomalous Green’s function and $\bm{g}=(g_{x},g_{y},g_{z})$ represents the attractive interaction between two electrons at an $s$-wave channel. To draw Fig. S1, we assume an attractive interaction at $p$-wave channel as usual
$g_{\nu}(\bm{k}-\bm{k}^{\prime})=g_{\nu}\cos(\theta-\theta^{\prime}),\;\cos\theta=k_{x}/k_{F},\;\sin\theta=k_{y}/k_{F}.$ (S12)
@@
To our knowledge, the two components proportional to $\hat{\xi}_{\bm{k}}$ at the second line in Eq. (S4) do not play any important role in stabilizing superconducting states. Namely, the presence of these components does not change $T_{c}$ at all. The first two components at the last line belong to odd-frequency symmetry class and make the superconducting state unstable [5]. A Zeeman field parallel to $\bm{d}$ and a SOI perpendicular to $\bm{d}$ generate such odd-frequency pairs from the pair potential. As a result, these components decrease $T_{c}$. The last component at the third line in Eq. (S4) belongs to even-frequency spin-singlet even-valley-parity symmetry class. This component stabilizes a superconducting state at high Zeeman fields. However, it appears only when the three vectors have a finite spin chiral product $\bm{\alpha}\times\bm{H}\cdot\,\bm{d}$.
@@
$\check{F}_{\parallel}(\bm{k},\omega_{n})=$ $-\frac{1}{2}\left[(\hat{X}_{+}^{-1}+\hat{X}_{-}^{-1})\bm{d}\cdot\hat{\bm{\sigma}}-(\hat{X}_{+}^{-1}-\hat{X}_{-}^{-1})d\right]i\hat{\sigma}_{y}\,(-i)\hat{\rho}_{y},$ (S13)
$\hat{X}_{\pm}$= $(\omega_{n}\pm iV)^{2}+(\hat{\xi}_{\bm{k}}\mp\alpha\,\hat{\rho}_{z})+d^{2}.$ (S14)
@@
$\check{F}_{\perp}(\bm{k},\omega_{n})=$ $-\check{Z}_{\perp}^{-1}\left[(\omega_{n}^{2}+\hat{\xi}_{\bm{k}}^{2}-\alpha^{2}+V^{2})\bm{d}\cdot\hat{\bm{\sigma}}-2i\hat{\xi}_{\bm{k}}\,\bm{d}\times\bm{V}\cdot\hat{\bm{\sigma}}\right$.
$\left.+2\omega_{n}\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}+2i\bm{\alpha}\times\bm{d}\cdot\bm{V}\hat{\rho}_{z}\right]i\hat{\sigma}_{y}(-i)\hat{\rho}_{y},$ (S15)
$\check{Z}_{\perp}=$ $\hat{\xi}_{\bm{k}}^{4}+2\hat{\xi}_{\bm{k}}^{2}(\omega_{n}^{2}-\alpha^{2}-V^{2})+(\omega_{n}^{2}+\alpha^{2}+V^{2})^{2}.$ (S16)